My first title

Arsenis Chrysikopoulos

March 8, 2021

1 My first section

My first paragraph

1.1 My first subsection

”Latex is a document preparation system and document markup language. LaTeX uses the Tex typesetting program for formating its output , and is

itself written in the Tex macro refers to the encoding or tagging conventions tha are used in LaTex documents. ”